OpenStudy (anonymous):

Find the price per unit p that produces the maximum profit P. C = 10x + 400 (Cost Function) p = 16 – 0.1√x (Demand function) $ ?

OpenStudy (anonymous):

So far I got worked out: P=R-C so, xp-(10x+400) =x(16-0.1sqrtx)-10x+400 =16-0.1sqrtx-10x+400? Am I doing this right? What do I do now?

OpenStudy (anonymous):

I really need help :(

OpenStudy (anonymous):

what is R? x*p?

OpenStudy (anonymous):

now you have profit as a function of x, maximize using calculus techniques (if you know them not sure)

OpenStudy (anonymous):

I believe he is saying R is revenue, that is quantity demanded times price.

OpenStudy (anonymous):

anyway what you are doing is wrong, it should be "xp-10x-400" as the minus is multiplying the whole term "10x-400"

OpenStudy (anonymous):

^^whoops she is right, I didn't catch that :P, but yeah take the function and maximize it, I'm not sure what level math you have had, or which econ this is though but thats how you can get the correct answer

OpenStudy (anonymous):

A quick edit, you will have found the maximum demand for the product, so you need to plug it back in to the demand function to get the price.

OpenStudy (anonymous):

I'm not sure how I'm doing it wrong.. i do have x*p p=16-0.1sqrtx... but I don't know what to do now..

OpenStudy (anonymous):

I don't know how to maximize :/

OpenStudy (anonymous):

This question is from my calculus homework. I'm taking it online and I have no help besides this study place, I need help walking me though the whole thing

OpenStudy (anonymous):

I have x(16-0.1sqrtx)-10x+400. what do I do with this?

OpenStudy (anonymous):

its x(16-0.1sqrtx)-10x-400 First multiply that first x out. It helps if you write it like such \[\frac{ x ^{1/2}}{ 10 }\] before multiplying out (:P you can add the exponents more easily)

OpenStudy (anonymous):

Next take d(Profit)/dx remember x is the quantity. Set it equal to zero. Because this is an econ question its safe to say this value will be the maximum, but usually you'd need to check it using a test (first derivative or second derivative).

OpenStudy (anonymous):

finally plug in your result into the demand function to find the p.

OpenStudy (anonymous):

so now it'S x^1/2/10=0?

OpenStudy (anonymous):

oh sorry I must have mislead you. I just told you write .1sqrtx in a different way to help you. you still need to multiply the whole thing out. go ahead and do it, tell me if you need help.

OpenStudy (anonymous):

ok, so i have 16-(x^1/2/10)-10x-400=0.. so x must equal..

OpenStudy (anonymous):

at least 0 right?

OpenStudy (anonymous):

nono hold it

OpenStudy (anonymous):

first take the expression and multiply it out, you get 16x- (1/10)x^3/2-10x-400

OpenStudy (anonymous):

where does the 1 come from?

OpenStudy (anonymous):

the 1/10x^3/2?

OpenStudy (anonymous):

uh im just writing x^3/2 divided by 10, its the same thing as .1x^3/2

OpenStudy (anonymous):

http://i.imgur.com/axwjYnZ.png

OpenStudy (anonymous):

ok.. so i start with 16x-(1/10)x^3/2-10x-400

OpenStudy (anonymous):

This is a plot of the profit curve with respect to quantity demanded, you can see the functions above it you are trying to find the max of this curve

OpenStudy (anonymous):

good, now take your equation you just typed, and take the derivative with respect to x and set it to 0, aka find the critical points (in this case one point because I know it is one)

OpenStudy (anonymous):

wow, I haven't even taken econ yet but I for sure can tell I will be confused when i take it lol

OpenStudy (anonymous):

don't worry too much about econ, lol the first year undergrad classes are extremely math easy. I'm in intermediate econ, and that is FINALLY when we use calc to maximize functions. If you are in High school, your econ is far easier than this problem

OpenStudy (anonymous):

if you are in college, like i said the first level classes are easy

OpenStudy (anonymous):

i'm a junior in college lol.. I still have a couple bac core classes before I can even start what i want to.. like this calculus, econ and physics.. anyhow, question at hand..

OpenStudy (anonymous):

i dunno how to take derivative of the first two terms because of the radical weird exponent stuff..

OpenStudy (anonymous):

i know the 400 drops off and the -10x is just -10

OpenStudy (anonymous):

good good, the first two terms you can take the derivative by using power rule. That is D[x^n]=n*x^(n-1)

OpenStudy (anonymous):

\[D[x^n]=n*x^{n-1}\]

OpenStudy (anonymous):

what do fractions look like as exponents come down in front from power rule :/

OpenStudy (anonymous):

the 16x also becomes 16, haha but i bet you knew that from what you said about the -10

OpenStudy (anonymous):

yeah.. the 16 is the easy one too lol

OpenStudy (anonymous):

fractions are the exact same, just follow the rule exactly, subtract one. 3/2-2/2=1/2

OpenStudy (anonymous):

so it just drops down to as 1/2x?

OpenStudy (anonymous):

so f' would be 16(1/2x)-10?

OpenStudy (anonymous):

not exactly, the 16x was a separate term (look above it was added) and the rule states you multiply by the original exponent, then drop it down

OpenStudy (anonymous):

so it would be \[6-\frac{ 3 }{ 2 }\frac{ p^{1/2} }{ 10 }\]

OpenStudy (anonymous):

wow, that just looks so complicated..

OpenStudy (anonymous):

^^ 16-10 simplifies to 6, and the 3/2 drops down while you subtract 1 from the exponent, as per the rule

OpenStudy (anonymous):

you can simplify further ofc, by multiplying the fraction out.

OpenStudy (anonymous):

\[6-\frac{ 3p^{1/2} }{ 20}\]

OpenStudy (anonymous):

set this equal to zero and solve for p. OH i just noticed this, i accidentally switched p and x, sorry about that, it is still x

OpenStudy (anonymous):

I'm bad at fractions and exponent.. everything else is fine for me..

OpenStudy (anonymous):

ok.. so now i set this to 0 and multiply it out?

OpenStudy (anonymous):

yeah, solve for x, (the p is x i accidentally switched it)

OpenStudy (anonymous):

after you solve for x, plug it into the demand function to find p.

OpenStudy (anonymous):

lol I actually used mathematica to model the problem and here is the answer for you to check. You should solve x=1600 and then plug it in and get the price to be 12 (as seen in the image)/ http://i.imgur.com/02b40VG.png

OpenStudy (anonymous):

right, gotcha.. geez, i was way off, i only got to 90 so far.. i was like geez... i need a big number to get this to equal 0

OpenStudy (anonymous):

do you need me to review the steps to solving it? I can do that for you too

OpenStudy (anonymous):

i was just trying to figure out the x.. so i just plug in the 1600 to the original function and it gives me 12 right?

OpenStudy (anonymous):

\[6=3\sqrt{x}/20\] \[120=3\sqrt{x}\] \[40=\sqrt{x}\] x=1600

OpenStudy (anonymous):

that last step is squaring both sides.

OpenStudy (anonymous):

oooooh.. so thats the easiest way to find x? square both sides?

OpenStudy (anonymous):

haha theres no surefire easy way, you just simplify the equation like you did in algebra. but in this case, you square both sides because the square root of x squared is x. so you need to square both sides to find x.

OpenStudy (anonymous):

right lol.. i hope to get the hang of this.. Thank you so very much for explaining this to me.. i would not have figured this out without you!

OpenStudy (anonymous):

just to review the complete answer is shown in the image. The quantity sold is 1600 the price is $12 and the profit is 2800.

OpenStudy (anonymous):

awesome! Thank you!!! I checked out the image :)